Which approximation was made when deriving the formula for finding bright spots?

a. tan(x) roughly equal to sin(x)

b. tan(x) roughly equal to cos(x)

c. sin(x) roughly equal to x

d. cos(x) roughly equal to x
Answer #1

In the derivation of formula for finding bright spots,

\sin x= \frac{PD}{d}

where PD is the path difference between the two light rays from the slits and d is the separation of the two slits

Also,

\tan x = \frac{y}{L}

where L is the distance from the slits to the screen and y is the distance of the first antinode from the central bright spot

The angle x is very small for this experiment, therefore the approximation applied is

tan(x) roughly equal to sin(x)